What is Sn of the geometric series with a1=3, an=-38, and n=8?
A) -178
B)-102
C)-140
D) -280

Answer by richwmiller(17219)

Tn = t * r^(n - 1)
-38 = 3 * r^(7)
r = -(38/3)^(1/7)
S=t*(1 - r^n)/(1 - r)
S=3*(1 - ( -(38/3)^(1/7))^8)/(1 - ( -(38/3)^(1/7)))
arithmetic
S = ½(a + an)n
S = ½(3 + -38)8=-140
This is probably meant to be an arithmetic series not GEOMETRIC